# Forecasting Response to Treatment with Global Deep Learning and Patient-Specific Pharmacokinetic Priors

### v6

@@ 2 Cohort
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to (Rubin-Falcone et al., 2022), the meal schedule used to generate simulated data was based on the Harrison-Benedict equation (Arthur Harris and Gano Benedict, 21919) with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
The ‘Dexcom’ continuous glucose monitor (CGM) option was used to obtain glucose readings at 5-minute intervals. The default basal-bolus controller provided in the simulation was used to administer insulin at the time of each meal, where the bolus amount is computed based on the current glucose level, the target glucose level, the patient’s correction factor, and the patient’s carbohydrate ratio. However, we adapt the controller to deliver basal doses once per hour, consistent with the OhioT1DM dataset. The resulting simulated data consists of CGM (in mg/dL) values, insulin rate (in units per minute), and carbohydrates (in grams), referred to as CHO. We extract the bolus insulin rate from the overall insulin rate as the difference in the continuous basal insulin rate, which is consistent hourly doses. Blood glucose time series for 6 simulated patients are shown in Fig. 7 in Appendix A.1.
Figure 3: Instantaneous medication administrations are represented as sparse variables in time series data. We propose a pharmacokinetic encoder to effectively capture time-dependent plasma drug concentration.
@@
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
Sparse exogenous features in the data include basal rate (in units per hour), temporary basal rate (units per hour), and bolus insulin (units), and CHO values. Bolus insulin consists of both ‘normal’ insulin, which is recorded as an instantaneous administration, and ‘square dual’, where the insulin dose is stretched over a defined period. ‘Square dual‘ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps. If multiple bolus administrations occurred within the same minute, the insulin units were summed. Basal insulin was included as hourly doses. For timestamps with overlapping basal rates, the last basal rate value was considered. When specified, the temporary basal rate superseded the basal rate.
@@
We use the specified train and test partitions provided in the dataset. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Missing values were forward filled to prevent data leakage. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1.
3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), based on a 10 hour history of blood glucose, carbohydrate, insulin basal, and insulin bolus values.
3.1 Pharmacokinetic Encoder
We propose a novel pharamacokinetic encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes. The encoder consists of a function, $C$, that takes in sparse exogenous medication variables and generates a concentration-time profile for a specified dose, $d$, using the following equation:
$C(t,d,\textbf{k}^{(i)})=\frac{d}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
where $t$ is the time and $\textbf{k}^{(i)}$ is the absorption rate constant specific to patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, $d$. Concentration profiles are generated under the assumption that all insulin administrations adhere to 100% bioavailability. As bioavailability equates to the area under the concentration curve, the scaled curve thus integrates to the specified dose, $d$.
Let $N$ be the number of subjects in the data. The absorption rate constant, $\textbf{k}_{\text{bolus}}\in[0,1]^{N\times 1}$ and $\textbf{k}_{\text{basal}}\in[0,1]^{N\times 1}$ are represented as network embedding weights that are initialized between values of 0 and 1 prior to model training. Initial $\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{bolus}}$ values are tuned as model hyperparameters. During model training, sparse basal and bolus features are passed into the pharmacokinetic encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder than replaces sparse medication features in the training window with the corresponding concentration curve. Non-medication features, such as CHO, are unaffected. For each training step, $\textbf{k}^{(i)}_{\text{bolus}}$ and $\textbf{k}^{(i)}_{\text{basal}}$ are updated for each patient, $i$, using gradient descent optimization.
For a single dose, $d$, of bolus insulin, $d_{\text{bolus}}$, or basal insulin, $d_{\text{basal}}$, administered at time, $t$, we obtain the concentration, $c$, at that time:
$c^{(i)}_{\text{bolus}}$ $=C(t,d^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)})$ (2)
$c^{(i)}_{\text{basal}}$ $=C(t,d^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)})$ (3)
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 3. The encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time, to perform vectorized operations that enable generating concentration curves for multiple doses in an input window in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps.
$\textbf{c}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}^{(i)}_{t},\textbf{k}^{(i)})$ (9)
Here, $f$ is the frequency of the data, $\textbf{d}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of doses, and $\textbf{c}^{(i)}\in\mathbb{R}^{1\times L}$ is the new plasma insulin concentration feature for patient, $i$.
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of concentration curves, c, other non-medication time series, x, and static features, s,
$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\textbf{c}^{(i)}_{t-L:t},\textbf{x}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (10)
In the proposed context, $\textbf{c}_{t-L:t}$ includes concentration curves for bolus insulin, $\textbf{c}_{\text{bolus}}$, and basal insulin $\textbf{c}_{\text{basal}}$, x includes CHO intake, and s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type, if applicable.
Figure 4: Hybrid Global-Local architecture. Global architectures with parameters shared across patients’ time series with local parameters from a patient-specific pharmacokinetic (Ph-K.) encoder. Sparse medication dose time series are passed into the pharmacokinetic encoder along with their respective $k_{a}$ parameters to generate concentration curves. Concentration curves are then passed into a deep learning model which learns relevant hidden states (hidden) and output parameters, $\hat{\theta}$.
3.2 Models
To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines. The implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries. More information on these models is provided in Appendix A.2.
Statistical Model
We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model.
Deep Learning Models
Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al. (2020); Challu et al. (2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al. (2014)), Long-Short Term Memory (LSTM; Sak et al. (2014)), Temporal Convolution Network with an MLP Decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)), and Temporal Fusion Transformer (TFT; Lim et al. (2021))
In contrast to previous efforts that focused on training models for individual subjects, or local models, we employ global architectures with parameters shared across subjects Rubin-Falcone et al. (2022); Xie and Wang (2020). For each dataset, models were trained using all patient data to generate one global model that was applied across patients. We showcase the efficacy of global architectures in producing improved forecasts by training local models for comparison against our global model. Moreover, apart from reducing computational and storage costs, global models could prove more applicable in real-world scenarios since they can be utilized for new patients with limited or no available training data.
Models were trained using Adam optimizer (Kingma and Ba, 2017). Optimal hyperparameters were selected based on a validation set with cross-validation. Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
Figure 5: The boxplots show average MAE computed across 8 trials for global (orange) and local (blue) models for OhioT1DM subjects. Models trained on all subject data, or global models, have lower MAE, on average, than models trained on individual subjects, or local models.
3.3 Baselines
We implement three baselines against which we compare our proposed hybrid global-local architexture and pharmacokinetic encoder approach:
Univariate signal
We train models with the univariate blood glucose signal.
Univariate signal with exogenous features
We train models with the univariate blood glucose signal, static features, and exogenous features, consisting of sparse basal and bolus insulin dose values.
Univariate signal with sum-total feature engineering
We train models with the univariate blood glucose signal, static features, and exogenous features, consisting of insulin dose values. The sum total feature preprocessing approach was used to transform sparse features into non-sparse features of cumulative dose across an input window, as shown in Fig. 3.
3.4 Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. In accordance with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models. Mathematical formulations for these metrics are provided in Appendix A.4. We evaluate model performance across all points within the specified time horizon. Furthermore, given the critical importance of healthcare outcomes related to blood glucose levels exceeding or falling below safe thresholds, we also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below certain thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
Our code implementation is open-source and can be found on Github. [^1] https://github.com/Nixtla/neuralforecast/tree/treat_module
@@ 4 Results
The pharmacokinetic encoder significantly improves forecasting accuracy for the top performing deep learning model by approximately 4.4% and 2.1%, for simulated and OhioT1DM datasets, respectively, as shown in Table 1. For the OhioT1DM dataset, the NHITS-Pharamacokinetic model has significantly smaller MAE than the NHITS-Sparse Exogenous model for forecasts evaluated at all values (t-test p-value: 4.44e-05) and critical values (t-test p-value: 1.6e-2). Fig. 6 demonstrates improved forecasting performance using the the NHITS-Pharamacokinetic model during hyperglycemic and hypoglycemic events for three randomly selected OhioT1DM subjects. Mean MAE and RMSE results with standard deviations computed across 8 trials are shown in Tables 7 and 8 in Appendix B.3.
We showcase the efficacy of our novel hybrid global-local model in producing more accuracy forecasts than models trained per patient. Global models consistently outperform local models, on average, for each OhioT1DM subject as shown in Table 4 and Fig. 9 in Appendix B.1. For the OhioT1DM dataset, the hybrid global-local pharmacokinetic model achieves a significantly lower (t-test p-value: 1.29e-2) MAE by 14.6% over patient-specific pharmacokinetic models. Fig. 5 shows that the global-local pharmacokinetic model has lower MAE, on average, across all OhioT1DM patients compared to the local models. Furthermore, the hybrid global-local pharmacokinetic model also achieves a significantly lower (t-test p-value: 3.83e-2) MAE by 9.2% over the patient-specific sparse exogenous models.
Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
@@
Baseline Exogenous Baseline Exogenous
ETS 10.040 13.531
MLP 9.734 8.653 12.189 9.887
LSTM 8.731 8.372 10.863 9.996
TCN 16.772 12.234 28.913 18.122
TFT 7.788 6.762 9.664 7.828
NHITS 8.333 7.282 10.184 8.353
TFT- Sum-Total 6.698 7.634
NHITS- Pharmacokinetic 7.027 7.917
TFT- Pharmacokinetic 6.466 7.514
@@
ETS 10.126 11.781
MLP 10.568 10.275 11.805 11.434
LSTM 10.604 9.606 12.999 11.144
TCN 11.011 11.090 12.693 13.212
TFT 9.605 9.515 10.686 10.555
NHITS 9.234 8.954 10.436 10.117
NHITS- Sum-Total 8.988 10.092
NHITS- Pharmacokinetic 8.769 10.007
TFT- Pharmacokinetic 9.492 10.484
Figure 6: Blood glucose forecasts of the final horizon point for three OhioT1DM subjects. The NHITS -Pharmacokinetic model (orange) produces noticeably improved forecasts of the ground truth (blue) compared to the NHITS model with sparse exogenous features (yellow).
@@ 5 Discussion
Our pharmacokinetic encoder enables deep learning models to achieve significantly lower forecasting errors. As such, models that only include sparse features may be under-leveraging the potential utility of exogenous information. Our pharmacokinetic encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings about unexpected treatment response, or helping to characterize patient-specific treatment effects in terms of drug absorption and elimination characteristics. The characterization of patient-specific pharmacokinetic parameters holds substantial potential for advancing the development of more effective treatments by tailoring medication to individual patients.
Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts. We examine whether the pharmacokinetic model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 10 in Appendix B.2. Our model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
Bolus insulin is quick-acting and often taken before meals compared to basal insulin, which is longer-acting. Our pharamcokinetic encoder learns $k_{a}$ values that align with this established knowledge. For OhioT1DM data, the average $k_{a_{bolus}}$ and $k_{a_{basal}}$ were 1.622 and 1.486, respectively. For our approach, larger $k_{a}$ values correspond to concentration curves with shorter duration, or steeper curves. The average $k_{a}$ values for bolus insulin were higher than those for basal insulin, on average, resulting in shorter durations for bolus insulin concentration profiles.
To evaluate whether the difference in $k_{a_{bolus}}$ and $k_{a_{basal}}$ values was statistically significant, we compute a one-sided paired t-test of model-inferred $k_{a}$ values for patients across the 8 trials. The alternative hypothesis claims that $k_{a}{{}_{bolus}}$ values are larger than $k_{a_{basal}}$. The test yielded a p-value of 0.043 indicating that the $k_{a}$ parameter for bolus insulin is significantly higher than that for basal insulin when considering a significance level of 0.05. Thus, our approach can learn the relative differences in pharamcokinetic characteristics between basal and bolus insulin.
@@
Limitations. We equate $\sigma$ to our $k_{a}$ parameter in the log-normal function and hold $\mu$ constant at 1. This simplification reduces the number of potential concentration curve shapes the model can learn. Tuning both $\sigma$ and $\mu$ would enable more flexible concentration curves with potentially more accurate duration and peak characteristics, which could further improve forecasts.
Future Work. Future work will consider adaptations to the concentrator log-normal function to enable more flexible concentration curves. Future work will also consider expanding our pharmacokinetic encoder with functions relevant to other medication administration methods, such as intravenous injections. Finally, another interesting line of research is determining how to leverage the advantages of the global forecasting model with patient-specific $k_{a}$ values for new patients.
@@
Appendix A First Appendix

### v7

@@ 4 Cohort
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. Blood glucose time series plots for 6 subjects are shown Fig. 7 in Appendix A.1. Basal and bolus dose distributions for 12 simuluted patients are shown in Fig. 9 in Appendix A.1.
@@
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The data includes basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1. Basal and bolus dose distributions for the 12 patients are shown in Fig. 10 in Appendix A.1.
More dataset information is in Appendix A.1.
@@ 5 Results
5.1 Results on Simulated Data
For the simulated dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous model for forecasts evaluated at all values (t-test p-value: 6.6e-6) as shown in as shown in Table 1. For evaluations across all values, the NHITS-PK model provides 3.5% improvement, on average, and up to 9.2% improvement for individual patients over the baseline NHITS sparse exogenous model. For evaluations across critical values, the NHITS-PK model provides 5.2% improvement, on average, and up to 16.4% improvement for individual patients over the baseline NHITS sparse exogenous model. Additionally, the NHITS-PK model outperforms the NHITS-ST model. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 6 in Appendix C.5.
5.2 Results on Real Data
Our hybrid global-local architecture produces models that significantly outperform local models by 15.8%, on average (t-test p-value: 1.7e-5), and up to 31.3% improvement for individual OhioT1DM patients. We show that the NHITS-PK hybrid global-local model outperforms NHITS-PK models developed for individual OhioT1DM patients in Fig. 12. Moreover, this improvement in forecast accuracy is consistent across all patients as shown in Fig. 5 and in Table 4 in Appendix C.1. Additionally, the NHITS-PK model with the hybrid global-local architecture also improves MAE by 14.8%, on average, for patients over the patient-specific NHITS sparse exogenous baseline models (t-test p-value: 1.7e-6).
For the OhioT1DM dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous baseline model for forecasts evaluated at all values (t-test p-value: 5.0e-3) and critical values (t-test p-value: 3.5e-3) as shown in Table 1. For evaluations across all values, The NHITS-PK model provides 1.3%  on average, and up to 3.5% improvement for individual patients over the baseline NHITS sparse exogenous model. For evaluations across critical values, the NHITS-PK provides 1.6%  on average, and up to 4.9% improvement for individual patients over the baseline NHITS sparse exogenous model. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 7 in Appendix C.5. Fig. 13 in Appendix C.2 shows model forecasts during hyperglycemic events for four randomly selected OhioT1DM patients.
Figure 5: The NHITS-PK hybrid global-local model (blue, solid) provides more accurate predictions than local NHITS-PK models (purple, dashed).
Figure 6: In addition to reducing forecasting error during hyperglycemic events, our NHITS-PK encoder enables improved lead time predictions of hypertensive events (blood glucose $\geq$ 180) for 9 of 12 OhioT1DM patients as measured over the 8 trials.
Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
@@ 5.2 Results on Real Data
Baseline Exog. Baseline Exog.
ETS 9.450 13.457
MLP 10.589 8.433 14.699 10.600
RNN 13.703 20.190 21.433 35.955
LSTM 13.557 10.321 20.537 13.550
TCN 12.096 10.109 17.709 12.660
NHITS 8.268 7.304 11.114 9.094
NHITS-ST 7.055 8.648
NHITS-PK 7.037 8.492
@@
ETS 9.022 10.028
MLP 11.065 11.120 12.630 12.698
RNN 11.325 11.795 13.627 13.552
LSTM 9.811 9.467 11.252 11.178
TCN 11.525 10.760 13.647 12.678
NHITS 9.060 8.873 10.337 10.137
NHITS-ST 8.958 10.131
NHITS-PK 8.758 9.965
@@ 6 Discussion
Hybrid Global-Local Architecture We demonstrate the efficacy of our novel hybrid global-local model, which produces forecasts with significantly smaller errors than patient-specific models across all patients. Our hybrid architecture not only improves accuracy but also enables global models to learn relevant patient-specific information, outperforming global model baselines.
A hybrid global-local model offers practical benefits beyond improved forecasting accuracy as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. The hybrid model can also be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
PK Encoder As a model-agnostic extension, our novel PK encoder significantly reduces forecasting errors in deep learning models for both simulated and real-world data. These improvements are not only statistically significant, but they have potential for positive real-world clinical impact. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts such events earlier, up to 3.6 minutes when compared to the NHITS baseline model, for individual OhioT1DM patients as shown in Fig. 6. This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations. Improving prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
Our results also highlight the importance of using domain knowledge to incorporate exogenous variables into models. Models with sparse features may underutilize valuable information, and feature-engineered representations not informed by domain knowledge, like the ‘sum total’ approach, can underperform compared to baselines. Moreover, Appendix C.4 shows that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset. As such, dose-dependent models, such as proposed in section 3.3, may be needed for more realistic data simulations and models.
Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing or drug selection. It can also be adapted for other medication administration types, such as oral or intravenous injections.
@@
Limitations. We equate $\sigma$ to our k parameter in the log-normal function and hold $\mu$ constant at 1. This simplification reduces the number of potential concentration curve shapes the model can learn. Tuning both $\sigma$ and $\mu$ could further improve forecasts by enabling more flexible concentration curves.
Future Work. Future work will explore fine-tuning global forecasting models with patient-specific parameters for new patients with limited historical data as well as modeling dose-dependent effects using more dynamic nonlinear PK parameter frameworks. Future work will include applying our approach to additional patient cohorts, such as those with T2DM, to further validate our findings.
@@
Appendix A Supplemental Methods
Data and Code Availability
We use open-source simulated and real-world data that is available to other researchers. The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository [Xie, 2018, Visentin et al., 2016]. The real-world data includes the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020]. Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu [2020].
Implementations of models used in our work were obtained from the open-source Neuralforecast [Olivares et al., 2022] and StatsForecast [Garza et al., 2022] libraries. We implemented our method using a copy of the Neuralforecast repository to leverage their model training framework. Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.